# How can Diffusion Models Evolve into Continual Generators?

### v2

@@ 4.4 Performance Comparison
Figure 2: Hyperparameter sensitivity analysis on the MNIST dataset, illustrating the effects of fine-tuning $\kappa$, $\lambda$, and $\eta$ on the metrics for MF and IMF.
Table 2: Ablations on each loss component.
Method MNIST
MF ($\downarrow$) IMF ($\downarrow$)
Effect of Removing Individual Loss Components
Baseline 67.60 48.79
$\mathcal{L}_{\text{IKC}}$ 57.92 45.46
$\mathcal{L}_{\text{IKC}}+\mathcal{L}_{\text{UKC}}$ 56.73 38.08
$\mathcal{L}_{\text{IKC}}+\mathcal{L}_{\text{LKC}}$ 50.70 43.79
$\mathcal{L}_{\text{IKC}}+\mathcal{L}_{\text{UKC}}+\mathcal{L}_{\text{LKC}}$ 48.97 40.08
4.5 Hyperparameter Fine-tuning
In this subsection, we conduct comprehensive hyperparameter tuning experiments associated with Eq. 1, using the MNIST dataset as the evaluation benchmark. Our sensitivity analysis (Fig. 2) highlights the pivotal role of three interrelated hyperparameters, $\kappa$, $\lambda$, and $\eta$, whose interplay governs CDG performance. The model consistency coefficient $\kappa$ achieves optimal MF and IMF scores (48.97 and 40.08) at $1\times 10^{-5}$. Increasing $\kappa$ to $1\times 10^{-3}$ significantly elevates MF (up to 293.36), suggesting overreliance on prior models, while reducing it to $1\times 10^{-7}$ causes a 18.2% drop in IMF, reflecting insufficient consistency. These trends show that moderate temporal coherence is essential for balancing adaptability and retention. The unconditional knowledge consistency weight $\lambda$ shows a similarly sensitive effect. At $1\times 10^{-5}$, IMF is minimized, whereas increasing it to $1\times 10^{-3}$ degrades IMF by 12.1%, indicating that overly strict constraints on the mean of unconditional sample embeddings may impede denoising flexibility. This underscores the importance of calibrated generative invariance. The label knowledge consistency coefficient $\eta$ also peaks at $1\times 10^{-5}$, effectively preserving cross-task semantic structure in the label space. These results confirm the utility of label-aware consistency in enhancing alignment in CDG and highlight the delicate yet critical balance among the components of our CCD framework. Each hyperparameter contributes uniquely to mitigating CF, and their joint calibration is essential for sustained CDG performance.
4.6 Ablation Studies
To systematically disentangle the contributions of each proposed loss component, we conduct controlled ablation studies on the MNIST benchmark (Tab. 2). The baseline model (MF: 67.60, IMF: 48.79) exhibits substantial forgetting, underscoring the vulnerability of unconstrained diffusion trajectories to knowledge drift across tasks. Introducing the inter-task knowledge consistency loss $\mathcal{L}_{IKC}$ significantly improves both metrics (MF: 57.92, IMF: 45.46), demonstrating the importance of Bregman-aligned model consistency in stabilizing forward diffusion dynamics. Adding the unconditional knowledge consistency loss $\mathcal{L}_{UKC}$ further improves performance (MF: 56.73, IMF: 38.08), indicating that aligning the mean of unconditional sample embeddings helps preserve reverse-time denoising trajectories. Substituting $\mathcal{L}_{UKC}$ with the label knowledge consistency loss $\mathcal{L}_{\text{LKC}}$ yields stronger gains in MF (50.70) but a smaller improvement in IMF (43.79), suggesting that label-space constraints are particularly effective for preserving shared semantics across tasks. When all three components are combined, the model achieves the best overall performance (MF: 48.97, IMF: 40.08), supporting our central hypothesis that multi-level consistency is critical for mitigating CF in CDG.
5 Conclusion
@@
This work introduces a principled framework for Continual Diffusion Generation (CDG), grounded in stochastic calculus, to tackle catastrophic forgetting (CF) in diffusion models. We characterize forgetting as misalignment in cross-task stochastic differential equation (SDE) dynamics and define three key consistency bounds, inter-task, unconditional mean, and label-level, to ensure stable knowledge retention. Based on these, we propose the Continual Consistency Diffusion (CCD) framework, which enforces multi-scale consistency via hierarchical loss functions that preserve the geometric and semantic integrity of generative trajectories. Experiments across diverse benchmarks show that CCD achieves state-of-the-art results in both Mean Fidelity (MF) and Incremental Mean Fidelity (IMF). By bridging static diffusion models and real-world streaming tasks, this work establishes CDG as a foundation for continual generative modeling in dynamic environments.
@@ A.1 Fundamental Concepts
By incorporating the Fokker-Planck equation and reverse-time dynamics in [3, 32], the corresponding reverse SDE can be formulated with an additional drift adjustment:
@@
The score estimator $\epsilon_{\theta}$ is employed to approximate the required $\nabla_{x_{t}}\log p_{t}(x_{t})$ for unconditional diffusion generation, and the optimization is achieved through the following loss function:
$\epsilon_{\theta}^{*}=\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{t}\Big{\{}\lambda(t)\,\mathbb{E}_{p(x_{0})}\mathbb{E}_{p(x_{t}|x_{0})}\big{[}\|\epsilon_{\theta}(x_{t},t)\;-\;\nabla_{x_{t}}\log p(x_{t}|x_{0})\|_{2}^{2}\big{]}\Big{\}}.$ (15)
@@
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})\;+\;\nabla_{x_{t}}\log p_{t}(y|x_{t})=\epsilon_{\theta}^{*}\;+\;\nabla_{x_{t}}\log p_{t}(y|x_{t}).$ (16)
A.2 Equivalence Between DDPM and SDE Models
In the Denoising Diffusion Probabilistic Model (DDPM) framework [9, 22], the data evolves through a forward process that progressively adds Gaussian noise to an initial sample $x_{0}$, resulting in a noisy sample $x_{t}$. The probability distribution of $x_{t}$ conditioned on $x_{0}$ can be written as:
@@ A.2 Equivalence Between DDPM and SDE Models
[…]
@@
A.3 Approximation of Score Functions and Noise Prediction
With the relationship between $x_{0}$ and $x_{t}$ established, we now turn to the connection between the score function in SDEs and the noise prediction model in DDPMs. Certainly, this is merely a rough justification. A more comprehensive and rigorous proof of the equivalence between the two can be found in [35]. In the SDE framework, the score function is defined as the gradient of the log-probability $p_{t}(x_{t}|x_{0})$ with respect to $x_{t}$:
@@ A.3 Approximation of Score Functions and Noise Prediction
$\epsilon_{\theta}(x_{t},t)$ $=\nabla_{x_{t}}\log p_{t}(x_{t})$ $\approx\mathbb{E}_{p(x_{0}),\,p(x_{t}|x_{0})}\bigl{[}\nabla_{x_{t}}\log p_{t}(x_{t}|x_{0})\bigr{]}$ $\approx-\frac{\varepsilon_{\theta}(x_{t},t)}{\bar{\beta}_{t}}.$ (28)
@@
Figure 3: Borrowing from [44], during multi-task optimization, the gradients will eventually converge. Therefore, in a streaming scenario, there exists a portion of shared knowledge, which ensures that $p_{t}(x_{t}|x_{0},y)=q_{t}(x_{t}|x_{0},y)$ and $\nabla_{x_{t}}\mu(x_{t},t)\approx\nabla_{x_{t}}\nu(x_{t},t)$. This shared knowledge is crucial for the diffusion generator’s ability to retain and transfer knowledge in continual generation.
@@ Appendix B Cross-Task Diffusion Evolution
From a Bayesian perspective, we analyze the forward transfer of diffusion models between two tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. The visual space distributions of these tasks can be represented as $p(x_{0})$ and $q(x_{0})$, where $x_{0}$ encompasses all visual samples within $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. The label spaces are expressed as $p(y)$ and $q(y)$, with $y$ representing all labels within $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. As illustrated in Fig. 3, the relationship between these tasks can be visualized through the joint distribution of $p(x_{0},y)$ and $q(x_{0},y)$. The task-specific distributions $p(x_{0})$ and $q(x_{0})$ share certain overlapping regions in the visual space, signifying common knowledge that can be transferred between the tasks. The key challenge is to leverage these shared regions while accounting for task-specific differences in visual and label spaces.
@@
$\nabla_{x_{t}}\log q_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t}|y)+\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\nabla_{x_{t}}\log\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right].$ (30)
Since $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ share some overlapping knowledge, we directly perform differential calculations on $x_{t}$. The key quantity of interest is the term $\nabla_{x_{t}}\log\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right]$, which can be expressed as the difference in noise terms between the tasks. And further combining with Eq. 16, we obtain:
@@
$\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}-\nabla_{x_{t}}\log\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right]+\nabla_{x_{t}}\log\frac{q_{t}(y|x_{t})}{p_{t}(y|x_{t})}$ (33)
$\leq$ $\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}+\nabla_{x_{t}}D_{KL}\Big{(}p_{t}(x_{0}|x_{t},y)\parallel q_{t}(x_{0}|x_{t},y)\Big{)}+\nabla_{x_{t}}\log\frac{q_{t}(y|x_{t})}{p_{t}(y|x_{t})}.$
@@
Applying Bayes’ theorem to decompose $p_{t}(x_{0}|x_{t},y)$, we obtain $p_{t}(x_{0}|x_{t},y)=\frac{p_{t}(x_{0}|x_{t})p_{t}(y|x_{0})}{p_{t}(y|x_{t})}$. Similarly, by performing an analogous decomposition for $q_{t}(x_{0}|x_{t},y)$, we arrive at:
@@
[…]
@@
$\quad+\underbrace{\nabla_{x_{t}}\int\frac{p_{t}(x_{0}|x_{t})p_{t}(y|x_{0})}{p_{t}(y|x_{t})}\log\frac{p_{t}(y|x_{0})}{q_{t}(y|x_{0})}dx_{0}}_{\text{Label Knowledge Consistency}}$
@@
$\nabla_{x_{t}}\!\int\frac{p_{t}(x_{0}|x_{t})\,p_{t}(y|x_{0})}{p_{t}(y|x_{t})}\log\frac{p_{t}(x_{0}|x_{t})}{q_{t}(x_{0}|x_{t})}\,dx_{0}$ $\;\propto\;\nabla_{x_{t}}\!\int p_{t}(x_{0}|x_{t})\,\log\frac{p_{t}(x_{0}|x_{t})}{q_{t}(x_{0}|x_{t})}\,dx_{0}$ (35)
@@
Hence, only the gradient of the KL divergence between the two posteriors $p_{t}(x_{0}|x_{t})$ and $q_{t}(x_{0}|x_{t})$ remains relevant to optimization, while the label-related factors play no direct role.
@@
$\nabla_{x_{t}}D_{KL}=\frac{(\mu(x_{t},t)-\nu(x_{t},t))}{\sigma_{t}^{2}}\quad(\nabla_{x_{t}}\mu(x_{t},t)-\nabla_{x_{t}}\nu(x_{t},t))\approx\delta\cdot\frac{\mu(x_{t},t)-\nu(x_{t},t)}{\sigma_{t}^{2}}.$ (36)
where $\delta$ represents a small perturbation determined based on the shared knowledge observed in Fig. 3.
@@
$\nabla_{x_{t}}D_{KL}\Big{(}p_{t}(x_{0}|x_{t})\parallel q_{t}(x_{0}|x_{t})\Big{)}=\frac{\bar{\alpha}_{t}^{2}}{\bar{\beta}_{t}^{2}}\Big{[}\mu_{\theta}(x_{t},t)-\nu_{\theta}(x_{t},t)\Big{]}.$ (37)
@@
Similarly, for the term representing Label Knowledge Consistency, the core focus is on directly describing the differences in the conditional distributions of labels $y$ given $x_{0}$. Here, the labels play a direct comparison role, and the label-related probability distributions directly impact the optimization objective. In contrast, $p_{t}(y|x_{t})$ does not affect the optimization objective related to $x_{0}$ and can be neglected to some extent. Therefore, this term can be transformed into:
@@
$p_{t}(x_{0}|x_{t})$ $=\frac{1}{\sqrt{(2\pi\sigma_{t}^{2})^{d}}}\times\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{0}^{p}-\mu(x_{t},t)\|^{2}\right),x_{0}^{p}\in\mathcal{T}_{i}.$ (39)
@@
$\nabla_{x_{t}}\log p_{t}(x_{0}|x_{t})$ $=\nabla_{x_{t}}\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{0}^{p}-\mu(x_{t},t)\|^{2}\right)=\frac{1}{\sigma_{t}^{2}}(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t).$ (40)
@@
$=\frac{\bar{\alpha}_{t}^{2}}{\bar{\beta}_{t}^{2}}\int p_{t}(x_{0}|x_{t})(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t)p_{t}(y|x_{0})\log\frac{p_{t}(y|x_{0})}{q_{t}(y|x_{0})}dx_{0}.$
@@
$(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t)\propto\frac{\bar{\beta}_{t}}{\bar{\alpha}_{t}}\|\varepsilon_{\theta}(x_{t}^{p},t)-\varepsilon\|_{2}^{2}=\frac{\bar{\beta}_{t}}{\bar{\alpha}_{t}}\|\beta_{t}\epsilon_{\theta}^{p}+\varepsilon\|_{2}^{2},$ (43)
where this step is based on the approximations derived in Eq. 28 and Equ. 43.
Ultimately, the term for Label Knowledge Consistency can be transformed into a form that is optimizable for SDE diffusion models:
@@
As shown above, the upper bound in Equ. 33 can be expressed as:
$\mathcal{L}_{UB}$ $=\kappa(\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p})+\lambda\,\mathcal{L}_{UKC}+\eta\,\mathcal{L}_{LKC}=\kappa\mathcal{L}_{IKC}+\lambda\,\mathcal{L}_{UKC}+\eta\,\mathcal{L}_{LKC},$ (45)
[…]

### v3

@@ Introduction
Figure 2: Hyperparameter sensitivity analysis on MNIST, illustrating the effects of fine-tuning $\kappa$, $\lambda$, and $\eta$.
Furthermore, to provide a more comprehensive evaluation, we include the various perceptual metrics in the Appendix.
Ablation Studies
In this subsection, we perform a sensitivity analysis on Eq. 1 using MNIST and an ER-5120 buffer strategy. As shown in Figure. 2, CDG performance is strongly influenced by three key hyperparameters: the model consistency coefficient $\kappa$, the unconditional generation consistency weight $\lambda$, and the prior knowledge consistency coefficient $\eta$. The optimal MF and IMF scores are observed when $\kappa=1\times 10^{-5}$. Increasing $\kappa$ leads to overreliance on prior models, while decreasing it weakens consistency and retention. Similarly, $\lambda=1\times 10^{-5}$ yields the best results, whereas higher values overly constrain the unconditional mean, limiting denoising flexibility. The prior knowledge consistency coefficient $\eta$ also peaks at $1\times 10^{-5}$, effectively preserving semantic structure across tasks. These results highlight the importance of carefully balancing temporal, generative, and semantic consistency. Each component plays a distinct role in mitigating GCF and their joint calibration is essential for the CCD performance.
Figure 3: Ablation studies on MNIST-5T (ER buffer 5120).
In addition, to further disentangle the individual contributions of each loss term, we also conduct controlled ablation studies on MNIST and an ER-5120 buffer strategy, summarized in Figure. 3. The baseline model (MF: 60.40, IMF: 135.19) suffers from severe forgetting and collapse, highlighting the instability of unconstrained diffusion across tasks. Adding $\mathcal{L}_{IKC}$ declines both MF and IMF, confirming the role of inter-task model consistency in stabilizing forward dynamics. Incorporating $\mathcal{L}_{UKC}$ further enhances performance by preserving reverse-time denoising through unconditional mean alignment. Replacing $\mathcal{L}_{UKC}$ with $\mathcal{L}_{PKC}$ yields lower MF but slightly higher IMF, indicating that label-space alignment is more effective for semantic retention. Combining all three losses achieves the best results (MF: 48.97, IMF: 40.08), validating that multi-level consistency is essential for mitigating forgetting in the CDG pipeline.
Visualization
To visualize how generative performance evolves in our CCD framework, we present a qualitative case study on CUB200-10T with an ER-5120 buffer strategy. Using the model trained through the final task, we generate $32\times 32$ images via a DDIM scheduler, sampling equally from Tasks 0, 4, and 8 (Figure. 3). Under standard ER replay, outputs collapse into near-pure noise, evidencing severe GCF. Introducing IKC and PKC losses progressively restores coarse contours, and the full CCD framework produces fully resolved images that permit unambiguous species identification. This stepwise reconstruction vividly illustrates CCD’s capacity to retain and consolidate knowledge across sequential tasks. For comprehensive visual examples, please see the Appendix.
Buffer Construction Strategies
Figure 4: Ablation study showcasing the generated images for tasks 0, 4, and 8 on the CUB200-10T benchmark using the model trained up to the final task.
Moreover, we observe that the first-in-first-out (FIFO) strategy in ER (Lopez-Paz and Ranzato 2017) introduces substantial randomness, often prioritizing redundant samples at the cost of diversity. Although CCD partially alleviates the resulting performance variance, it compromises consistency across tasks. To address this limitation, we introduce a Hierarchical Diversity Buffer (HDB) architecture.
HDB: To counteract the instability caused by FIFO-based sample replacement, HDB partitions memory into two components: a Candidate Pool for temporary intake and an Elite Repository for long-term storage of diverse exemplars. When the Candidate Pool reaches capacity, samples are assessed via similarity scoring using an exponential decay kernel, $S_{ij}=\exp(-\|\mathbf{h}_{i}-\mathbf{h}_{j}\|^{2}/2\alpha^{2})$, where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ denote prior regression embeddings from PKC. Samples with low average similarity (i.e., high diversity) are promoted to the Elite Repository. This progressive filtering maximizes coverage of the underlying data distribution while avoiding redundancy, leading to more stable and effective rehearsal-based training.
As shown in Table 2, HDB further improves CCD performance on coarse-grained datasets. However, we also observe certain adverse effects on fine-grained datasets. This limitation largely stems from the non-discriminative nature of intermediate outputs in diffusion models, making it challenging for HDB, and most CL methods, to effectively determine which samples should be retained. This insight points to a promising direction for future work: enhancing the discriminative quality of intermediate diffusion representations.
[…]
MNIST-5T 50.05 23.96 47.55 22.28 44.27 22.12
OxfordPets-5T 287.21 271.76 288.65 279.51 286.64 275.56
CIFAR100-10T 102.34 93.60 93.36 85.08 90.89 80.70
Flowers102-10T 90.31 97.37 81.43 81.41 66.38 80.01
CUB200-10T 140.60 100.52 142.66 99.20 140.77 99.77
Table 2: Performance of our CCD optimization framework across different datasets under varying HDB buffer sizes.
Conclusion
@@
This work presents a principled framework for our standardized CDG pipeline, grounded in stochastic calculus, to mitigate GCF in diffusion models. We conceptualize forgetting as a misalignment in cross-task SDE dynamics and introduce three key consistency constraints, inter-task, unconditional mean, and prior space, to promote stable knowledge retention. Building on these insights, we propose the CCD framework, which enforces multi-level consistency through hierarchical loss functions that preserve both the geometric and semantic integrity of generative trajectories. Experiments on diverse benchmarks demonstrate that CCD achieves state-of-the-art performance in both MF and IMF. By bridging static diffusion modeling with real-world streaming tasks, CCD lays a solid foundation for continual generation tasks.
@@ Appendix A Preliminaries in SDE Diffusion Models
By incorporating the Fokker-Planck equation and reverse-time dynamics in (Anderson 1982; Song et al. 2021), the corresponding reverse SDE can be formulated with an additional drift adjustment:
@@
The score estimator $\epsilon_{\theta}$ is employed to approximate the required $\nabla_{x_{t}}\log p_{t}(x_{t})$ for unconditional diffusion generation, and the optimization is achieved through the following loss:
$\epsilon_{\theta}^{*}=$ $\mathop{\arg\min}\limits_{\theta}\,\mathbb{E}_{t}\bigg{\{}\lambda(t)\,\mathbb{E}_{p(x_{0})}\big{[}\mathbb{E}_{p(x_{t}|x_{0})}\|\epsilon_{\theta}(x_{t},t)$
$-\nabla_{x_{t}}\log p(x_{t}|x_{0})\|_{2}^{2}\big{]}\bigg{\}}$ (15)
@@
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)$ $=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})$
$=\epsilon_{\theta}^{*}+\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (16)
Equivalence Between DDPM and SDE Models
In the Denoising Diffusion Probabilistic Model (DDPM) framework (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021), the data evolves through a forward process that progressively adds Gaussian noise to an initial sample $x_{0}$, resulting in a noisy sample $x_{t}$. The probability distribution of $x_{t}$ conditioned on $x_{0}$ can be written as:
@@
$x_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)ds}x_{0}+\int_{0}^{t}\sqrt{\beta(s)}e^{-\frac{1}{2}\int_{s}^{t}\beta(u)du}dW_{s}.$ (21)
By choosing $\bar{\alpha}(t)=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)ds}$ and $\bar{\sigma}^{2}(t)=1-\bar{\alpha}^{2}(t)$, we can write:
@@
Score Approximation & Noise Prediction
With the relationship between $x_{0}$ and $x_{t}$ established, we now turn to the connection between the score function in SDEs and the noise prediction model in DDPMs. Certainly, this is merely a rough justification. A more comprehensive and rigorous proof of the equivalence between the two can be found in (Vincent 2011). In the SDE framework, the score function is defined as the gradient of the log-probability $p_{t}(x_{t}|x_{0})$ with respect to $x_{t}$:
@@
Figure 5: Borrowing from (Yu et al. 2020), during multi-task optimization, the gradients will eventually converge. Therefore, in a streaming scenario, there exists a portion of shared knowledge, which ensures that $p_{t}(x_{t}|x_{0},y)=q_{t}(x_{t}|x_{0},y)$ and $\nabla_{x_{t}}\mu(x_{t},t)\approx\nabla_{x_{t}}\nu(x_{t},t)$. This shared knowledge is crucial for the diffusion generator’s ability to retain and transfer knowledge.
@@
$\epsilon_{\theta}(x_{t},t)$ $=\nabla_{x_{t}}\log p_{t}(x_{t})$ (28)
$\approx\mathbb{E}_{p(x_{0}),\,p(x_{t}|x_{0})}\bigl{[}\nabla_{x_{t}}\log p_{t}(x_{t}|x_{0})\bigr{]}$
$\approx-\frac{\varepsilon_{\theta}(x_{t},t)}{\bar{\beta}_{t}}.$
@@ Appendix B Cross-Task Diffusion Evolution
From a Bayesian perspective, we analyze the forward transfer of diffusion models between two tasks $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. The visual space distributions of these tasks can be represented as $p(x_{0})$ and $q(x_{0})$, where $x_{0}$ encompasses all visual samples within $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. The prior spaces are expressed as $p(y)$ and $q(y)$, with $y$ representing all labels within $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$. As illustrated in Figure. 5, the relationship between these tasks can be visualized through the joint distribution of $p(x_{0},y)$ and $q(x_{0},y)$. The task-specific distributions $p(x_{0})$ and $q(x_{0})$ share certain overlapping regions in the visual space, signifying common knowledge that can be transferred between the tasks. The key challenge is to leverage these shared regions while accounting for task-specific differences.
@@
$\nabla_{x_{t}}\log q_{t}(x_{t}|y)$ $=\nabla_{x_{t}}\log p_{t}(x_{t}|y)$ (30)
$+\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\nabla_{x_{t}}\log\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right].$
Since $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ share some overlapping knowledge, we directly perform differential calculations on $x_{t}$. The key quantity of interest is the term $\nabla_{x_{t}}\log\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right]$, which can be expressed as the difference in noise terms between the tasks. And combining with Eq. 16, we obtain:
@@
$\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}-\nabla_{x_{t}}\log\mathbb{E}_{p_{t}(x_{0}|x_{t},y)}\left[\frac{q_{t}(x_{0}|y)}{p_{t}(x_{0}|y)}\right]$ (33)
$\quad+\nabla_{x_{t}}\log\frac{q_{t}(y|x_{t})}{p_{t}(y|x_{t})}$
$\leq$ $\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}+\nabla_{x_{t}}D_{KL}\Big{(}p_{t}(x_{0}|x_{t},y)\parallel q_{t}(x_{0}|x_{t},y)\Big{)}$
$\quad+\nabla_{x_{t}}\log\frac{q_{t}(y|x_{t})}{p_{t}(y|x_{t})}.$
@@
Applying Bayes’ theorem to decompose $p_{t}(x_{0}|x_{t},y)$, we obtain $p_{t}(x_{0}|x_{t},y)=\frac{p_{t}(x_{0}|x_{t})p_{t}(y|x_{0})}{p_{t}(y|x_{t})}$. By performing an analogous decomposition for $q_{t}(x_{0}|x_{t},y)$, we arrive at:
@@
$=\nabla_{x_{t}}\int p_{t}(x_{0}|x_{t},y)\Bigg{[}\log\frac{p_{t}(x_{0}|x_{t})}{q_{t}(x_{0}|x_{t})}+\log\frac{p_{t}(y|x_{0})}{q_{t}(y|x_{0})}$
$\quad+\log\frac{q_{t}(y|x_{t})}{p_{t}(y|x_{t})}\Bigg{]}dx_{0}$
@@
$\quad+\underbrace{\nabla_{x_{t}}\int\frac{p_{t}(x_{0}|x_{t})p_{t}(y|x_{0})}{p_{t}(y|x_{t})}\log\frac{p_{t}(y|x_{0})}{q_{t}(y|x_{0})}dx_{0}}_{\text{Prior Knowledge Consistency}}$
@@
$\nabla_{x_{t}}\!\int\frac{p_{t}(x_{0}|x_{t})\,p_{t}(y|x_{0})}{p_{t}(y|x_{t})}\log\frac{p_{t}(x_{0}|x_{t})}{q_{t}(x_{0}|x_{t})}\,dx_{0}$ (35)
$=\nabla_{x_{t}}\!\int p_{t}(x_{0}|x_{t})\,\log\frac{p_{t}(x_{0}|x_{t})}{q_{t}(x_{0}|x_{t})}\,dx_{0}$
@@
Hence, only the gradient of the KL divergence between the two posteriors $p_{t}(x_{0}|x_{t})$ and $q_{t}(x_{0}|x_{t})$ remains relevant to optimization, while the label-related factors play no role.
@@
$\nabla_{x_{t}}D_{KL}$ $=\frac{(\mu(x_{t},t)-\nu(x_{t},t))}{\sigma_{t}^{2}}\big{(}\nabla_{x_{t}}\mu(x_{t},t)-\nabla_{x_{t}}\nu(x_{t},t)\big{)}$ (36)
$\approx\delta\cdot\frac{\mu(x_{t},t)-\nu(x_{t},t)}{\sigma_{t}^{2}}.$
where $\delta$ represents a small perturbation determined based on the shared knowledge observed in Figure. 5.
@@
$\nabla_{x_{t}}D_{KL}\Big{(}p_{t}(x_{0}|x_{t})\parallel q_{t}(x_{0}|x_{t})\Big{)}$ (37)
$=\frac{\bar{\alpha}_{t}^{2}}{\bar{\beta}_{t}^{2}}\times\Big{[}\mu_{\theta}(x_{t},t)-\nu_{\theta}(x_{t},t)\Big{]}.$
@@
Similarly, for the term representing Prior Knowledge Consistency, the core focus is on directly describing the differences in the conditional distributions of labels $y$ given $x_{0}$. Here, the labels play a direct comparison role, and the label-related probability distributions directly impact the optimization objective. In contrast, $p_{t}(y|x_{t})$ does not affect the optimization objective related to $x_{0}$ and can be neglected to some extent. Therefore, this term can be transformed into:
@@
$p_{t}(x_{0}|x_{t})$ $=\frac{1}{\sqrt{(2\pi\sigma_{t}^{2})^{d}}}\times\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{0}^{p}-\mu(x_{t},t)\|^{2}\right)$ (39)
$\quad\text{where }x_{0}^{p}\in\mathcal{T}_{i}.$
@@
$\nabla_{x_{t}}\log p_{t}(x_{0}|x_{t})$ $=\nabla_{x_{t}}\left(-\frac{1}{2\sigma_{t}^{2}}\|x_{0}^{p}-\mu(x_{t},t)\|^{2}\right)$ (40)
$=\frac{1}{\sigma_{t}^{2}}(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t).$
@@
$=\frac{\bar{\alpha}_{t}^{2}}{\bar{\beta}_{t}^{2}}\int p_{t}(x_{0}|x_{t})(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t)p_{t}(y|x_{0})$
$\quad\times\log\frac{p_{t}(y|x_{0})}{q_{t}(y|x_{0})}dx_{0}.$
@@
$(x_{0}^{p}-\mu(x_{t},t))\nabla_{x_{t}}\mu(x_{t},t)$ $\propto\frac{\bar{\beta}_{t}}{\bar{\alpha}_{t}}\|\varepsilon_{\theta}(x_{t}^{p},t)-\varepsilon\|_{2}^{2}$ (43)
$=\frac{\bar{\beta}_{t}}{\bar{\alpha}_{t}}\|\beta_{t}\epsilon_{\theta}^{p}+\varepsilon\|_{2}^{2}.$
where this step is based on the approximations derived in Eq. 28 and Eq. 43.
Ultimately, the term for Prior Knowledge Consistency can be transformed into a form that is optimizable for SDE diffusion models:
@@
As shown above, the upper bound in Eq. 33 can be expressed as:
$\mathcal{L}_{UB}$ $=\kappa(\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p})+\lambda\,\mathcal{L}_{UKC}+\eta\,\mathcal{L}_{PKC}$ (45)
$=\kappa\mathcal{L}_{IKC}+\lambda\,\mathcal{L}_{UKC}+\eta\,\mathcal{L}_{PKC}.$
$\left\{\begin{aligned} &\mathcal{L}_{IKC}=\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p},\\ &\mathcal{L}_{UKC}=\frac{\bar{\alpha}_{t}^{2}}{\bar{\beta}_{t}^{2}}\left[\mu_{\theta}(x_{t},t)-\nu_{\theta}(x_{t},t)\right],\\ &\mathcal{L}_{PKC}=\frac{\bar{\alpha}_{t}}{\bar{\beta}_{t}}\mathbb{E}_{p_{t}(x_{0}|x_{t})}D_{KL}(p_{t}(y|x_{0})||q_{t}(y|x_{0})),\end{aligned}\right$. (46)